OpenStudy (anonymous):

I dont understand complex Fractions it says Which expression is equivalent to the complex fraction but i dont know how to figure it out! I just need someone to walk me through it please!!!! A. 2/X B. X/2 C.2x/X^2 + 4 D. 2x/X+2

OpenStudy (anonymous):

oh shoot forgot attachment

OpenStudy (australopithecus):

set \[1 = \frac{x+2}{x+2}\] then subtract the bottom fractions to make one fraction then just flip the bottom fraction and multiply it by the top fraction

OpenStudy (australopithecus):

so \[\frac{\frac{x}{x+2}}{\frac{x+2}{x+2}-\frac{x}{x+2}}\]

OpenStudy (australopithecus):

since they have the same denominator you can combine them to \[\frac{\frac{x}{x+2}}{\frac{x+2-x}{x+2}} = \frac{\frac{x}{x+2}}{\frac{2}{x+2}}\]

OpenStudy (anonymous):

So the answer would be 2x ------------ X^2 +4?

OpenStudy (australopithecus):

\[\frac{\frac{x}{x+2}}{\frac{2}{x+2}} = \frac{x}{x+2}(\frac{x+2}{2})\]

OpenStudy (australopithecus):

follow my steps and try to understand I dont want you asking this question again on here lol

OpenStudy (australopithecus):

if you have any questions ask

OpenStudy (australopithecus):

When ever dividing you flip the fraction

OpenStudy (australopithecus):

on the bottom and multiply

OpenStudy (australopithecus):

do you not know how to multiply fractions? it is really simple

OpenStudy (australopithecus):

do you understand why 1+x/1+x = 1

OpenStudy (anonymous):

Yes they cancel each other out

OpenStudy (australopithecus):

any quantity that is divided by itself is equal to 1 1 + x is just a number

OpenStudy (australopithecus):

it is just manipulation of the numbers and it is important to know :)

OpenStudy (australopithecus):

so do you know what to do next?

OpenStudy (australopithecus):

as far as multiplying fractions

OpenStudy (anonymous):

No i dont

OpenStudy (australopithecus):



OpenStudy (australopithecus):

 now can you do it